# Galaxy Groups in the presence of Cosmological Constant: Re-Mapping the close-by galaxies

### v1

@@ V Multipole expansion
The two examples for one-dominant and two-dominant galaxies are simple cases for a general configuration of galaxies. The gravitational potential $\Phi(r)$ inside the group or cluster is the solution of the Poisson equation:
$\Delta\Phi=4\pi G\rho-\Lambda c^{2}.$ (31)
For a static case, the integration of this equation yields the following:
@@
$M(r)=\int\rho(r^{\prime})d^{3}r^{\prime},$
${\bf p}=\int\rho(r^{\prime})\left(\hat{\bf r}-{\bf r}^{\prime}\right)d^{3}r^{\prime}=0,$
@@
The expansion is around the CoM in the origin, therefore ${\bf p}=0$  [^1] Notice that the dipole, octuple and and all the multipoles with odd powers in ${\bf r}_{i}$ defined in the centre of mass system (or defined in any other system as odd powers in $\vec{r}_{i}-\vec{r}_{\text{CoM}}$) of any system should be equal to zero, due to the parity invariance of gravitational theories.
The normalization by the total mass gives:
$\sigma_{\text{Quad}}=\frac{1}{M}\sum_{ij}Q_{ij}\hat{r}^{i}\hat{r}^{j},$ (36)
in terms of the basis components of the quadrupole tensor. Normalizing again by the monopole zero acceleration surface ($\bar{r}_{0}$), we obtain a dimensionless parameter that quantifies the validity condition for multipole expansion in our formalism: $\tilde{\sigma}_{\text{Quad}}=\sigma_{\text{Quad}}/\bar{r}_{0}^{2}\ll 1$.
@@
where $A$ is a function of the angular variables. Instering the solution to Eq. (37) gives:
@@
VI Conclusions
@@
We extended the definitions of characteristic surfaces for galaxy groups to include the effect of dark energy. In this general framework, the interplay between the attraction of matter and dark-matter distribution and the repulsion force of dark energy gives a natural definition for the local groups. In our approach, we model the dwarf galaxies as test particles around the giant galaxies. While Tully (2015) defines the groups of galaxy without the presence of dark energy, we extend the definitions to the boundaries of galaxy groups with the local effect of the cosmological constant $\Lambda$.
@@
For the monopole case, where there is one giant galaxy, there is a clear radius where the total force is zero, namely the zero gravity radius. By considering the system to obey the virial theorem, one can shows that this radius is also the upper radius for the virilizied system. For a galaxy pair, the total acceleration for the monopole case is generalized to the Zero Radial Acceleration Surface (ZRAS) that obeys Eq. (5). The turnaround surface is obtained by setting the kinetic energy of dwarf galaxies to be positive. In this case, a potential surface that obeys Eq. (6) is obtained and gives a boundary for the group.
@@
With these two simple examples, we map the galaxies in the local Universe based on the data from CosmicFlow. We consider different giant galaxies and galaxy pairs, and we show that the close-by groups are separated. The dynamics of these groups can be determined by the outside and inside the boundary: Inside the boundary, the system will become closer to a spherically symmetric solution in the future (the pair will fall into the center of mass) and the groups will continue to distance themselves apart from each other. We generalize these solution using multipole expansion, and we show that choosing the correct center of mass of system yields an elegant general description: The system could be determine by the monopole and the qudrupole moments (since the dipole is zero) and the boundaries could be determine by the interplay of gravitational attraction and dark energy.
@@ Acknowledgements.
We thank Noam Libeskind, Till Sawala, Horst Stocker and Jenny Wagner for illuminating discussions.

### v3

@@ V Multipole expansion
The two examples for one-dominant and two-dominant galaxies are simple cases for a general configuration of galaxies. The gravitational potential $\Phi(r)$ inside the group or cluster is the solution of the Poisson equation: $\Delta\Phi=4\pi G\rho-\Lambda c^{2}$. For a static case, the integration of this equation yields the following:
@@
$M(r)=\int\rho(r^{\prime})d^{3}r^{\prime},\quad{\bf p}=\int\rho(r^{\prime})\left(\hat{\bf r}-{\bf r}^{\prime}\right)d^{3}r^{\prime}=0,$
@@
The expansion is around the CoM in the origin, therefore ${\bf p}=0$  [^1] Notice that the dipole, octupole and and all the multipoles with odd powers in ${\bf r}_{i}$ defined in the centre of mass system (or defined in any other system as odd powers in ${\bf r}_{i}-{\bf r}_{\text{CoM}}$) of any system should be equal to zero, due to the parity invariance of gravitational theories.. The normalization by the total mass gives:
$\sigma_{\text{Quad}}(\theta,\phi)=\frac{1}{M}\sum_{ij}Q_{ij}\hat{r}^{i}\hat{r}^{j},$ (27)
in terms of the basis components of the quadrupole tensor. Normalizing again by the zero acceleration surface ($\bar{r}_{0}$) for the monopole case, we obtain a dimensionless parameter that quantifies the validity condition for multipole expansion in our formalism: $\tilde{\sigma}_{\text{Quad}}=\sigma_{\text{Quad}}/\bar{r}_{0}^{2}\ll 1$.
@@
where $A$ is a function of the angular variables. Inserting the solution to Eq. (28) gives:
@@
VI Estimated Local Group Masses with $\Lambda$
To show the importance of $\Lambda$ in a galaxy group, we estimate its contribution to the masses of these groups. We modeled the groups to be dominating by one or two giant galaxies $(M1,M2)$ and taking $N$ remote galaxies of similar average mass $m$. Based on our derivation, we use their detailed locations and velocities from the CoM for these two galaxies. From Eq. (8), the kinetic energy average is:
$\left\langle K\right\rangle=\frac{3}{2}mN\sigma_{v}^{2},$ (33)
where $\sigma_{v}$ is the one-dimensional velocity dispersion of the system, and the factor of three accounts for the total dispersion under the assumption of isotropy.
Since the measurement is along the line-of-sight velocities, one has to subtract the Hubble flow via: $\mathbf{v}_{\text{LoS}}=\mathbf{v}_{i}+H_{0}\mathbf{r}_{i}$, where $\mathbf{v}_{\text{LoS}}$ is the measured line-of-sight velocity and $r_{i}$ is the vector distance to the individual galaxy, and $v_{i}$ is the line-of-sight component of the peculiar motion. We assume that self-gravity of the outer galaxy groups is negligible compared to the external influence inside the group. The average gravitational potential is:
$\left\langle U_{G}\right\rangle=-GmN\left(\frac{M_{1}}{r_{1G}}+\frac{M_{2}}{r_{2G}}\right),$ (34)
where we define the length $r_{G}=N^{2}\sum_{i}1/r_{i0}$ that runs between the dwarf galaxies with respect to the giant ones. The average dark energy potential is:
$\left\langle U_{\Lambda}\right\rangle=\frac{1}{2}mN\Omega_{\Lambda}H_{0}^{2}\sigma_{r}^{2}.$ (35)
Altogether, the virial theorem yields:
$3\sigma_{v}^{2}+\frac{1}{2}\Omega_{\Lambda}H_{0}^{2}\sigma_{r}^{2}=GM\left(\frac{\gamma}{r_{1G}}+\frac{1-\gamma}{r_{2G}}\right),$ (36)
where $\gamma=M_{1}/M$. For a system with only one giant galaxy, $M_{2}=0$, and then we consider the center to be around $M_{1}$. For a given ratio of the giants, it is possible to infer the total mass from the dispersions of velocities and the distances. For a comparison Diaz et al. (2014); Hartl and Strigari (2022) uses $\gamma=M_{31}/M_{MW}=2.3$ and obtains $\left(2.5\pm 0.4\right)\cdot 10^{12}\,M_{\odot}$. The contribution of $\Lambda$ increases the predicted mass by the ratio:
$M_{\text{VT}+\Lambda}=M_{\text{VT}}\left[1+\frac{\Omega_{\Lambda}}{6}\left(\frac{H_{0}\sigma_{r}}{\sigma_{v}}\right)^{2}\right],$ (37)
where $M_{VT}$ is the estimated mass considering the Newtonian interaction and $M_{VT+\Lambda}$ is the estimated mass with $\Lambda$.
We assume a Gaussian prior on: $H_{0}=67.4\pm 0.5\,km/s/Mpc$, $\Omega_{\Lambda}=0.685\pm 0.007$ and 5% error on the velocities. The enclosed mass of the LG is estimated to be: $M_{LG}^{(VT)}=\left(2.33\pm 0.07\right)\cdot 10^{12}M_{\odot}$. With $\Lambda$, the mass increases to $M_{LG}^{(VT+\Lambda)}=\left(2.47\pm 0.08\right)\cdot 10^{12}M_{\odot}$. Fig. (4) shows the posterior distribution for the estimated masses with and without $\Lambda$. The mass is shifted by $\sim 5-10\%$ which is within the errors of the estimated masses. However, with future measurements, especially around the turnaround surface, it will be possible to increase the precision and the effect of $\Lambda$ will be seen as a stronger one.
Figure 4: The posterior distribution for the LG total mass using the virial theorem with and without $\Lambda$. $\Lambda$ creates a repulsion force and to compensate for this additional interaction, the virialized mass of the system has to be higher.
VII Conclusions
@@
The present work extends the definitions of characteristic surfaces for the galaxy groups to include the effect of dark energy. In this general framework, the interplay between the attraction of matter and dark-matter distribution and the repulsion force of dark energy gives a natural definition for the local groups. Within this approach, we model the dwarf galaxies as test particles around the giant galaxies. In earlier studies, galaxies were grouped on the basis of the observed distribution of the considered galaxies, assuming only Newtonian interaction, and subsequently found the masses enclosed in the groups. In the present work, however, we emphasize the importance of dark energy inside the galaxy groups and in addition we extend the definitions of the boundaries of galaxy groups by including the local effect of the cosmological constant $\Lambda$.
@@
For the monopole case, where there is one giant galaxy, there is a clear radius where the total force is zero, namely the zero gravity radius. By considering the system to obey the virial theorem, one can shows that this radius is also the upper radius for the virialized system. For a galaxy pair, the total acceleration for the monopole case is generalized to the Zero Radial Acceleration Surface (ZRAS) that obeys Eq. (3). The turnaround surface is obtained by setting the kinetic energy of dwarf galaxies to be positive. In this case, a potential surface that obeys Eq. (4) is obtained and gives a boundary for the group.
@@
With these two simple examples, we map the galaxies in the local Universe based on the data from CosmicFlow. We consider different giant galaxies and galaxy pairs, and we show that the close-by groups are separated. The dynamics of these groups can be determined by the outside and inside the boundary: Inside the boundary, the system will become closer to a spherically symmetric solution in the future (the pair will fall into the center of mass) and the groups will continue to distance themselves apart from each other. We generalize these solution using multipole expansion, and we show that choosing the correct center of mass of system yields an elegant general description: The system could be determine by the monopole and the quadrupole moments (since the dipole is zero) and the boundaries could be determine by the interplay of gravitational attraction and dark energy.
By invoking the powerful Virial Theorem  [^2] The Virial Theorem has previously been used in astrophysics in Milgrom (2014) for the modified gravitational theory, MOND; see also Milgrom (1994) for a general derivation of the theorem for action-governed theories, and as well for some historical references therein on many applications of the theorem., we estimate the LG mass, assuming that dwarf galaxies around MW and M31 have a very small mass. We show that the estimated LG mass increases by 5 to 10% to be $\left(2.47\pm 0.08\right)\cdot 10^{12}M_{\odot}$.
The result is consistent with other independent measurements of the LG mass. Specifically, it aligns with values reported in the literature, which also suggest $\sim[2,3]\cdot 10^{12}M_{\odot}$ based on fitting the motions of dwarf galaxies to the Hubble flow (Peirani and de Freitas Pacheco, 2006; Peirani and Pacheco, 2008; Karachentsev et al., 2009; Peñarrubia et al., 2014; Teerikorpi and Chernin, 2010; Del Popolo and Chan, 2022). Additionally, considering the Milky Way and Andromeda (M31) as a binary system with the Large Magellanic Cloud acting as a perturbation Benisty (2024), a mass of $\left(2.3\pm 0.7\right)\cdot 10^{12}M_{\odot}$ has been derived.
We compared our total mass estimates with the sum of the individual galaxy masses. Various studies provide values for the separate masses of each galaxy: $M_{MW}\in[0.9,1.3]\cdot 10^{12}M_{\odot}$ and $M_{31}\in[1.,2.]\cdot 10^{12},M_{\odot}$ (Wang et al., 2020; Sawala et al., 2023a). The total mass of both galaxies falls within the range $M_{MW}+M_{31}\in[1.9,3.3]\cdot 10^{12}M_{\odot}$. The LG mass should at least be within this range, though higher values might arise due to the presence of undetected dark matter around the galaxies. The total mass estimated from the virial theorem (including $\Lambda$) is consistent with this range, suggesting that additional dark matter between the galaxies may not be necessary. However, this remains an open question, as some studies, such as Ref. Sawala et al. (2023b), propose additional dark matter between the galaxies based on simulations.
The whole work shows the importance of invoking the effect of $\Lambda$, which has previously been ignored in studies of the local universe. Moreover, we show that the definition of groups naturally comes from the interplay between the effect of $\Lambda$ with other constituents of matter and their interaction, and therefore should be considered in future studies, even within close-by scales.
@@ Acknowledgements.
We are grateful to Noam Libeskind, Mordehai Milgrom, Paolo Salucci, Horst Stocker, and in particular Jenny Wagner for the illuminating discussions. DB is supported by a Minerva Fellowship of the Minerva Stiftung Gesellschaft fuer die Forschung mbH.